# Density Fluctuations in the Intracluster Medium: An Attempt to Constrain Viscosity with Cosmological Simulations

### v1

@@ 3.4 Thermodynamical structure
For a better understanding of the gas dynamics of the clusters, in this section, we study the thermodynamical structure of the different clusters. For a better comparison with observations, we focus only on the inner regions of the cluster (within $R_{\mathrm{2500c}}$), as we did in previous sections. To visualize the thermodynamical structure of the clusters, Fig. 10 shows a 2D histogram of the line-of-sight velocity ($v_{\mathrm{LoS}}$) as a function of the temperature, color-coded by the emissivity. For the calculation of the emissivity, we assume it to be $\propto\rho^{2}\sqrt{T}$, expected for thermal bremsstrahlung (Sarazin, 1986). We also overplot the mass distribution of the particles as a contour for the different $v_{\mathrm{LoS}}$ and temperatures.
Figure 10: 2D histogram of the $v_{\mathrm{LoS}}$ as a function of the temperature, color-coded by emissivity. The contours indicate the mass distribution of the particles within the cluster. From left to right: MHD only, MHD with $1/3$ of Spitzer viscosity and full Spitzer viscosity at redshift zero. From top to bottom: clusters g55, g16 and g63.
@@
It is important to note that cluster g16 exhibits a narrower mass-weighted temperature distribution in the three configurations. The majority of the mass is concentrated in a narrow range of temperatures, meaning that this cluster is closer to be isothermal for the three runs than the other two clusters. This is similar to the case analysed by Zhuravleva et al. (2019), where they assumed an isothermal fluid of $T\sim 9.28\times 10^{7}$ K for Coma. The narrower the temperature distribution is, the narrower is the range of viscosities as well (see equation 9). This would imply a sharper cutoff in the turbulent cascade, leading to a steeper 3D amplitude density fluctuations. Although cluster g16 is not perfectly isothermal, this steeper 3D amplitude density fluctuations can be seen in the middle panel of Fig. 7 by comparing the non-viscous and the viscous case. The maximum produced by the choice of $f_{\mathrm{cut}}$ might affect the results, however the steeper spectrum can also be seen in Fig. 13 for different values of $f_{\mathrm{cut}}$. The difference in slope is not so prominent as the one suggested by Zhuravleva et al. (2019), however it indicates that their criterion to constrain viscosity might be applicable only to isothermal galaxy clusters. The consequence of the narrower temperature distribution can also be seen in the middle panel of Fig. 15, where the dynamical range of the spectrum reaches smaller scales in the non-viscous than in the viscous case.
@@ 4 Conclusions
In this work we performed a set of cosmological simulations using the code OpenGadget3 of three different galaxy clusters with three different values of viscosity to quantify the amount of viscosity in the ICM. First, we compared the results of these simulations at redshift zero in order to understand the effect that viscosity has in the ICM. Then we tried to constrain the amount of viscosity in the ICM by comparing with X-ray observations of density fluctuations. Our key conclusions are:
  • Although the overall morphology remains the same, by visual inspection one can identify morphological differences produced by viscosity due the suppression of instabilities at small scales. The runs with more viscosity show a larger amount of small clumps that have not been disrupted by instabilities; a more filamentary structure produced by the gas stripped from infalling structures towards the center of the cluster; and a more concentrated gas in the denser regions, rather than the more mixed and homogeneous gas seen in the non-viscous cases.
@@
  • Using a fixed value of $f_{\mathrm{cut}}$ to divide the bulk and high-density gas works well for inviscid simulations of galaxy clusters. However, the results with full viscosity (where the density PDF is broader) are slightly dependent on the choice of $f_{\mathrm{cut}}$. Therefore, a more accurate method to separate bulk and high-density gas should be investigated in the future to avoid the dependence on the shape of the density PDF.
  • The density and velocity fluctuations depend linearly and both increase with the distance to the center. However, in all three clusters the slope of the velocity-density fluctuations decreases the more viscous the medium is. This linear relation is translated into a density to velocity fluctuations ratio of the order of unity for distances up to $R_{\mathrm{200c}}$.
  • The runs with viscosity tend to have larger amplitude of density fluctuations as a function of the scale. However, the density fluctuations obtained from our simulations are consistent with observations, even in the case with full viscosity. This is due to the large range of temperature distribution. In isothermal clusters, the slope of the amplitude of density fluctuations is affected by viscosity. This would mean that the method suggested in Zhuravleva et al. (2019) to constrain the amount of viscosity in the ICM by measuring the slope of the amplitude of density fluctuations is only applicable to isothermal clusters.
@@ Acknowledgments
TM would like to thank Irina Zhuravleva and Eugene Churazov for the intense discussions which motivated this work. KD and TM acknowledge support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. UPS is supported by the Simons Foundation through a Flatiron Research Fellowship (FRF) a the Center for Computational Astrophysics (CCA). The CCA is supported by the Simons Foundation. MV is supported by the Italian Research Center on High Performance Computing, Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2, within the activities of Spoke 3, Astrophysics and Cosmos Observations, and by the INFN Indark Grant. DVP has been supported by the Agencia Estatal de Investigación Española (AEI; grant PID2022-138855NB-C33), by the Ministerio de Ciencia e Innovación (MCIN) within the Plan de Recuperación, Transformación y Resiliencia del Gobierno de España through the project ASFAE/2022/001, with funding from European Union NextGenerationEU (PRTR-C17.I1), by the Generalitat Valenciana (grant CIPROM/2022/49), and by Universitat de València through an Atracció de Talent fellowship. This work has been supported by the Munich Excellence Cluster Origins funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2094 390783311. The simulations were performed at the Computational Center for Particle and Astrophysics (C2PAP).
@@ Appendix A Dependence on $f_{\mathrm{cut}}$
Viscosity broadens the density PDF due to the lack of mixing. As a result, it becomes difficult to split the bulk and the high-density gas. The median value of density is displaced towards larger values of density, producing that the threshold calculated using equation 18 cannot split bulk and high-density gas accurately. In the most extreme cases, this threshold lies outside the range of densities for certain shells of the cluster, as can be seen in Fig. 11 for cluster g16 for $f_{\mathrm{cut}}=2.5$.
@@
This signifies that some high-density clumps are not removed, leading to a maximum in the density fluctuations spectrum, as shown in Fig. 7. This can be solved by reducing the value of $f_{\mathrm{cut}}$ in equation 18 to lower values than $2.5$. By doing this, we can successfully split bulk and high-density gas, as can be seen in the upper panel of Fig. 12, where we used $f_{\mathrm{cut}}=0.5$. As a consequence, the maximum observed in the amplitude of the density fluctuations is reduced (see Fig. 13). However, if we use $f_{\mathrm{cut}}=0.5$ in other clusters, the split of bulk and high-density gas is not properly done (see bottom panel of Fig. 12 to observe the effect in cluster g55). This value of $f_{\mathrm{cut}}$ would consider part of the bulk gas as high-density gas and would remove more gas than only high-density clumps. For consistency we use the same value of $f_{\mathrm{cut}}$ in all our clusters, choosing the value suggested in Zhuravleva et al. (2012) of $f_{\mathrm{cut}}=2.5$ and acknowledging the spurious effects that it can have in the amplitude of density fluctuations.
@@ Appendix B Density profile
If we look at the density radial profile of each one of the clusters (Fig. 14) we do not observe any clear trend among the cases with different viscosities. This shows how at large scales the viscosity effects do not significantly affect the cluster morphology. We need to study the density fluctuations ($\sim 5$%-$10$% of the total density (Churazov et al., 2012; Sanders & Fabian, 2012)) to see bigger differences between the cases with and without viscosity. When doing a radial profile, the mean value of density is calculated for each shell, removing this way any effect due to viscosity. Fig. 3 shows how, even though the density distribution is broader with viscosity, the mean value remains more or less the same, leading to very similar radial density profiles.
@@ Appendix C Velocity power spectrum
We can make use of the code vortex-p to interpolate the particles into a grid and compute the energy power spectrum for each cluster and for each amount of viscosity in each case. Fig. 15 shows the power spectrum normalised to Kolmogorov for each cluster within $R_{\mathrm{2500c}}$. In all cases the slope is steeper than Kolmogorov, however, the slope is the same in each cluster regardless the amount of viscosity. At larger scales the runs with viscosity appear to be more energetic, although at intermediate and small scales the viscosity runs become less energetic (except for the cluster g16).

### v2

@@ 3.4 Thermodynamical structure
For a better understanding of the gas dynamics of the clusters, in this section, we study the thermodynamical structure of the different clusters. For a better comparison with observations, we focus only on the inner regions of the cluster (within $R_{\mathrm{2500c}}$), as we did in previous sections. To visualize the thermodynamical structure of the clusters, Fig. 10 shows 2D histograms of the line-of-sight velocity ($v_{\mathrm{LoS}}$) as a function of the temperature, color-coded by the emissivity. For the calculation of the emissivity, we assume it to be $\propto\rho^{2}\sqrt{T}$, expected for thermal bremsstrahlung (Sarazin, 1986). We also overplot the mass distribution of the particles as a contour for the different $v_{\mathrm{LoS}}$ and temperatures.
Figure 10: 2D histograms of the $v_{\mathrm{LoS}}$ as a function of the temperature, color-coded by emissivity. The contours indicate the mass distribution of the particles within the cluster. From left to right: MHD only, MHD with $1/3$ of Spitzer viscosity and full Spitzer viscosity at redshift zero. From top to bottom: clusters g55, g16 and g63.
@@
It is important to note that cluster g16 exhibits a narrower mass-weighted temperature distribution in the three configurations. The majority of the mass is concentrated in a narrow range of temperatures, meaning that this cluster is closer to being isothermal for the three runs than the other two clusters. This is similar to the case analysed by Zhuravleva et al. (2019), where they assumed an isothermal fluid of $T\sim 9.28\times 10^{7}$ K for Coma. The narrower the temperature distribution is, the narrower is the range of viscosities as well (see equation 9). This would imply a sharper cutoff in the turbulent cascade, leading to a steeper 3D amplitude density fluctuations. Although cluster g16 is not perfectly isothermal, these steeper 3D amplitude density fluctuations can be seen in the middle panel of Fig. 7 by comparing the non-viscous and the viscous case. The maximum produced by the choice of $f_{\mathrm{cut}}$ might affect the results, however the steeper spectrum can also be seen in Fig. 13 for different values of $f_{\mathrm{cut}}$. The difference in slope is not as prominent as the one suggested by Zhuravleva et al. (2019), however, it indicates that their criterion to constrain viscosity might be applicable only to isothermal galaxy clusters. The consequence of the narrower temperature distribution can also be seen in the middle panel of Fig. 15, where the dynamical range of the spectrum reaches smaller scales in the non-viscous than in the viscous case.
@@ 4 Conclusions
In this work we performed a set of cosmological simulations using the code OpenGadget3 of three different galaxy clusters, each with three different values of viscosity to quantify the effect of viscosity in the ICM. First, we compared the results of these simulations at redshift zero in order to understand the effect that viscosity has in the ICM. Then we tried to constrain the amount of viscosity in the ICM by comparing with X-ray observations of density fluctuations. Our key conclusions are:
  • Although the overall morphology remains the same, by visual inspection one can identify morphological differences produced by viscosity due the suppression of instabilities on small scales. The runs with more viscosity show a larger amount of small clumps that have not been disrupted by instabilities; a more filamentary structure produced by the gas stripped from infalling structures towards the center of the cluster; and more gas concentrated in the denser regions, rather than the more mixed and homogeneous gas seen in the non-viscous cases.
@@
  • Using a fixed value of $f_{\mathrm{cut}}$, used to divide the bulk and high-density gas, works well for inviscid simulations of galaxy clusters. However, the results with full viscosity (where the density PDF is broader) are slightly dependent on the choice of $f_{\mathrm{cut}}$. Therefore, a more accurate method to separate bulk and high-density gas should be investigated in the future to avoid the dependence on the shape of the density PDF.
  • The density and velocity fluctuations are directly proportional to one another and both increase with the distance from the center. However, in all three clusters the slope of the relationship between velocity and density fluctuations decreases with increasing viscosity. This linear relation is translated into a density to velocity fluctuations ratio of the order of unity for distances up to $R_{\mathrm{200c}}$.
  • The runs with viscosity tend to have larger amplitude of density fluctuations as a function of the scale. However, the density fluctuations obtained from our simulations are consistent with observations, even in the case with full viscosity. This is due to the large range of temperature distribution. In isothermal clusters, the slope of the amplitude of density fluctuations is affected by viscosity. This means that the method suggested in Zhuravleva et al. (2019) to constrain the amount of viscosity in the ICM by measuring the slope of the amplitude of density fluctuations is only applicable to isothermal clusters.
@@ Acknowledgments
TM would like to thank Irina Zhuravleva and Eugene Churazov for the intense discussions which motivated this work. The authors also want to thank the referee for their very useful comments. KD and TM acknowledge support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. UPS is supported by the Simons Foundation through a Flatiron Research Fellowship (FRF) a the Center for Computational Astrophysics (CCA). The CCA is supported by the Simons Foundation. MV is supported by the Italian Research Center on High Performance Computing, Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2, within the activities of Spoke 3, Astrophysics and Cosmos Observations, and by the INFN Indark Grant. DVP has been supported by the Agencia Estatal de Investigación Española (AEI; grant PID2022-138855NB-C33), by the Ministerio de Ciencia e Innovación (MCIN) within the Plan de Recuperación, Transformación y Resiliencia del Gobierno de España through the project ASFAE/2022/001, with funding from European Union NextGenerationEU (PRTR-C17.I1), by the Generalitat Valenciana (grant CIPROM/2022/49), and by Universitat de València through an Atracció de Talent fellowship. This work has been supported by the Munich Excellence Cluster Origins funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2094 390783311. The simulations were performed at the Computational Center for Particle and Astrophysics (C2PAP).
@@ Appendix A Dependence on $f_{\mathrm{cut}}$
Viscosity broadens the density PDF due to the lack of mixing. As a result, it becomes difficult to define a dividing line between the bulk and the high-density gas. The median value of density is displaced towards larger values of density, causing the threshold defined in equation 14 to be a poor choice for the division between bulk and high-density gas. In the most extreme cases, this threshold lies outside the range of densities for certain shells of the cluster, as can be seen in Fig. 11 for cluster g16 for $f_{\mathrm{cut}}=2.5$.
@@
This signifies that some high-density clumps are not removed, leading to the maximum in the density fluctuations spectrum, as shown in Fig. 7. This can be solved by reducing the value of $f_{\mathrm{cut}}$ in equation 14 to lower values than $2.5$. By doing this, we can successfully split bulk and high-density gas, as can be seen in the upper panel of Fig. 12, where we used $f_{\mathrm{cut}}=0.5$. As a consequence, the maximum observed in the amplitude of the density fluctuations is reduced (see Fig. 13). However, if we use $f_{\mathrm{cut}}=0.5$ in other clusters, the split of bulk and high-density gas is not properly done (see bottom panel of Fig. 12 to observe the effect in cluster g55). This value of $f_{\mathrm{cut}}$ would consider part of the bulk gas as high-density gas and would remove more gas than only high-density clumps. For consistency we use the same value of $f_{\mathrm{cut}}$ in all our clusters, choosing the value suggested in Zhuravleva et al. (2012) of $f_{\mathrm{cut}}=2.5$ and acknowledging the spurious effects that it can have in the amplitude of density fluctuations.
@@ Appendix B Density profile
If we look at the radial density profile of each one of the clusters (Fig. 14) we do not observe any clear trend among the cases with different viscosities. This shows how at large scales the viscosity effects do not significantly affect the cluster morphology. We need to study the density fluctuations ($\sim 5$%-$10$% of the total density (Churazov et al., 2012; Sanders & Fabian, 2012)) to see bigger differences between the cases with and without viscosity. When making a radial profile, using the mean density for each shell hides the main impact of the viscosity, which is seen in the fluctuations about the mean. Fig. 3 shows how, even though the density distribution is broader with viscosity, the mean value remains more or less the same, leading to very similar radial density profiles.
@@ Appendix C Velocity power spectrum
We can make use of the code vortex-p to interpolate the particles into a grid and compute the velocity power spectrum for each cluster and for each amount of viscosity in each case. Fig. 15 shows the power spectrum normalised to Kolmogorov for each cluster within $R_{\mathrm{2500c}}$. In all cases the slope is steeper than Kolmogorov, however, the slope is the same in each cluster regardless the amount of viscosity. At larger scales the runs with viscosity appear to be more energetic, although at intermediate and small scales the viscosity runs become less energetic (except for the cluster g16).